# Improving Medical Multi-modal Contrastive Learning with Expert Annotations

### v1

@@ 1 Introduction
Pretraining foundation models on multi-modal data – particularly leveraging the relationships between text and images – has proven to be a robust strategy for generating versatile embeddings (Radford et al., 2021; Jia et al., 2021). These embeddings enhance the efficacy in several downstream tasks, from image generation to advanced vision-language integration (Liu et al., 2023; Ramesh et al., 2021; Singh et al., 2022). Central to this approach is the employment of a contrastive learning (CL) loss objective (Oord et al., 2018; Chen et al., 2020; Zbontar et al., 2021), where models are trained to align positive pairs (e.g., an image and its corresponding caption) while diversifying negative ones. A significant hurdle in this approach is the necessity of vast datasets, often comprising several millions of data points, for competitive results. Models such as CLIP (Radford et al., 2021) have been trained on internet-scale datasets, estimated to encompass hundreds of millions of image-text pairs (Cherti et al., 2023; Xu et al., 2023a). Acquiring datasets of this magnitude poses substantial challenges in specialized fields that require expert knowledge for data collection, processing and annotation. The medical imaging domain exemplifies these difficulties, where acquiring even a single data point, such as a chest X-ray, involves complex processes requiring expertise and significant resources. Moreover, the procurement of such data for machine learning research is further complicated by ethical considerations, patient privacy concerns and the need for extensive de-identification procedures.
@@
This has led to the prevalent use of foundation models, initialized with weights from models trained on extensive internet-scale datasets, for tasks in the medical domain (Zhang et al., 2022; Wang et al., 2022; Huang et al., 2021; Krishnan et al., 2022). However, the areas of interest within medical images are often nuanced and require expert knowledge to interpret, rendering them indistinguishable to a general-purpose model. In 0(a), we investigate the embeddings generated by a CLIP model – initially pretrained on internet data – using samples from the Open-I dataset(Demner-Fushman et al., 2016), which includes X-rays and corresponding radiology reports. We categorize the samples into subgroups based on the primary abnormality identified in each report, such as ‘normal’, cardiomegaly, atelectasis and opacity. A histogram of the cosine similarities between embeddings from different groups indicates a high degree of similarity, with values approaching 1. This could lead to potential challenges in downstream zero-shot inference tasks, which rely on the spatial segregation of embeddings from different groups (Radford et al., 2021). Typically, continual pretraining on medically relevant data is employed to enhance the model’s ability to differentiate between various abnormalities.
@@
((a)) Histogram of Cosine Similarities Among Subgroups. The model exhibits high cosine similarity among embeddings within the text (red) and image (blue) modalities, regardless of the differences among subgroups. This underscores the model’s challenge in capturing the subtleties inherent in medical data.
@@
((b)) Modality Gap. Despite the CLIP contrastive loss aiming at closely aligning image and text embeddings within a shared space, the modalities remain segregated into distinct regions.
@@
Recent studies (Liang et al., 2022; Goel et al., 2022; Oh et al., 2023; Zhang et al., 2024; Tschannen et al., 2023) have identified a “modality gap” in multi-modal contrastive representation learning, where the embeddings from different modalities (e.g., images and text) fall in distinct regions in the shared embedding space. This separation, which arises from factors such as initial model weights and the objectives of contrastive learning (Liang et al., 2022; Zhang et al., 2024), leads to the “cone effect” where embeddings of each modality are restricted to a narrow region of the embedding hypersphere. In 0(b), we illustrate this within the medical domain with a 2D UMAP(McInnes et al., 2018) projection of image and text embeddings. This example highlights how embeddings from the same modality but different semantic groups, such as X-ray images of varying abnormalities, cluster closely together. This makes it difficult for a model to distinguish between semantically different images, undermining its performance in medical image analysis.
@@
We investigate the potential of integrating expert annotations, specifically radiologist eye-gaze heatmaps, to alleviate these issues. Processing the eye-gaze data from radiologists(Karargyris et al., 2021) provides us with heatmaps indicative of the radiologist’s attention across different regions of the X-ray images. This heatmap reflects areas of clinical interest aligned with details present in radiology reports. We posit that this could help capture nuanced visual cues in the X-rays and therefore pairing it with reports can enrich the CLIP training data with high-quality positive pairs. Due to the scarcity of such expert annotated data, we employ the mixup strategy, a data augmentation technique which has been effective in both supervised (Zhang et al., 2017; Verma et al., 2019; Han et al., 2022) and contrastive learning(Verma et al., 2021; Oh et al., 2023), to create additional synthetic samples.
@@ 2 Related Work
Modality Gap: Liang et al. (2022) pinpoint the origins of the modality gap to the nuances of model initialization and the objectives of contrastive learning, underscoring its impact on downstream tasks and fairness. Oh et al. (2023) highlight poor uniformity and alignment in CLIP’s embeddings and propose a finetuning method for robust representations. Zhang et al. (2024) explore the geometry of this embedding space, and provide both theoretical and empirical insights on the nature of this geometry. Subsequent research has produced methods to mitigate the modality gap through diverse and creative approaches (Goel et al., 2022; Tschannen et al., 2023; Zhang et al., 2023b; Gu et al., 2023; Nukrai et al., 2022).
Improving Contrastive Learning: Several methods have improved upon the CLIP objective by introducing auxilliary losses, e.g., SLIP (Mu et al., 2022) uses SimCLR (Chen et al., 2020) loss, M3MAE (Geng et al., ; Weers et al., 2023) augment the Masked Autoencoder (He et al., 2022) reconstruction loss, FLIP (Li et al., 2023) randomly masks out input images to improve scaling, DACL (Verma et al., 2021) proposes a domain agnostic mixup strategy, SILC (Naeem et al., 2023) uses self-distillation, Mo et al. (2024) utilized specialist captions to generate pseudo labels for unpaired images, Zhang et al. (2023a) propose Multi-task Paired Masking with Alignment to improve cross-modal interaction. Similarly there have been works that have identified the need to make the CLIP model focus on sub-regions in order to enhance its utility and downstream performance, GLoRIA (Huang et al., 2021) considers the loss from local regions from within the image and reports, Alpha-CLIP (Sun et al., 2023) uses the alpha channel to guide the CLIP model to focus on different regions of the image and generate the masks for all the images in the corpus using an image segmentation pipeline and TIER (Palepu & Beam, 2023) uses a regularization term to improve the local focus of the model.
Multi-modal Contrastive Learning in Medical Imaging: Zhang et al. (2022) demonstrated enhanced downstream performance by jointly using chest X-ray and report pairing for training a contrastive learning model. This was further improved by Huang et al. (2021), by exploiting local and global features from both modalities; Wang et al. (2022) and You et al. (2023) achieved impressive results by using a Swin Tiny model as the image encoder and by adding modifications to the contrastive loss. Several other works have developed similar contrastive learning foundation models while utilizing the biomedical image texts to achieve impressive results (Huang et al., 2023; Xu et al., 2023b; Moor et al., 2023; Wu et al., 2023; Sowrirajan et al., 2010; Tu et al., 2024). Karargyris et al. (2021) and Bigolin Lanfredi et al. (2022) augment a subset of the MIMIC-CXR (Johnson et al., 2019b) samples with high quality eye-tracking and verbal transcripts from several radiologists. van Sonsbeek et al. (2023) and Wang et al. (2024) utilize the heatmaps from eye-gaze to improve image classification.
@@ 3.1 Background
Central to CLIP’s effectiveness is the InfoNCE loss (Oord et al., 2018), a mechanism engineered to optimize the similarity measures between corresponding (positive) pairs and to minimize those among non-corresponding (negative) pairs. The formulation of the CLIP loss objective is as follows:
@@
Theoretical results on CL indicate the concepts of alignment and uniformity as critical for the quality of embeddings (Wang & Isola, 2020; Wang & Liu, 2021). Alignment focuses on reducing the distance between positive pairs while uniformity seeks to evenly distribute the embeddings across the unit hypersphere, preventing extreme clustering that could impair the model’s generalizability and discriminative capabilities. The alignment and uniformity can be defined formally as follows (Oh et al., 2023):
@@
Figure 3: Comparing eCLIP with $m^{2}$-mixup(Oh et al., 2023). (left) Standard CLIP showing image-text positive pairs $(v_{i},t_{i})$ (solid line), while the other image embeddings serve as negative pairs (dashed line). (center) the $m^{2}$-mixup creates negative pairs $(v^{\lambda}_{j},t_{i})$ via interpolation between embeddings along the geodesic. (right) eCLIP adds expert image embedding, $v^{E}_{i}$, in addition to $v_{i}$ for text $t_{i}$, forming additional positive and negative pairs
High intra-modal similarity, e.g. between two images as seen in Fig. 1, can inadvertently enhance similarity among negative pairs, inflating the denominator of the loss function in Equation (1), and, consequently, hurting the model’s ability to differentiate between positive and negative pairs during training. A conventional method to counter this involves incorporating hard negative pairs, a strategy (Oh et al., 2023) employ by mixing embeddings from different modalities. While effective, this cross-modal mixup may obscure the semantic clarity of embeddings. As an alternative we propose increasing the dataset with additional positive pairs that exhibit minimal semantic overlap by integrating expert annotations. Fig. 3 compares the positive and negative pair creation in eCLIP with traditional CLIP and $m^{2}$-mixup.
@@ 3.2 Introducing Expert Annotations to CLIP
8:        $\lambda\sim\text{Beta}(\alpha,\alpha)$
9:        $I_{i}^{E}\leftarrow\text{HeatmapProcessor}(I_{i},E_{i})$
10:        $I_{i}^{\lambda}\leftarrow I_{i}\lambda+I_{i}^{E}(1-\lambda)$
11:        $v_{i}^{\lambda}\leftarrow f(I_{i}^{\lambda})$
12:     end if
13:     if $E_{i}$ is entirely ones then
14:        $I_{i}^{\text{R}}\leftarrow\text{HeatmapProcessor}(I_{i},E_{i})$
15:        $\mathcal{L}_{\text{priming}}\leftarrow\mathcal{L}_{\text{priming}}+(I_{i}-I_{i}^{\text{R}})^{2}$
16:        $n_{\text{p}}\leftarrow n_{\text{p}}+1$
17:     end if
18:  end for
@@
However, the size of the expert annotated data is orders of magnitude smaller than the data available for CLIP training. To effectively leverage the scarce expert-annotated data, we implement mixup augmentation (Zhang et al., 2017). As illustrated in Fig. 2 (left), this involves blending an original image $I_{i}$ with its expert version $I_{i}^{E}$ to create $I_{i}^{\lambda}=\lambda I_{i}+(1-\lambda)I_{i}^{E}$, where $\lambda\sim\text{Beta}(\alpha,\alpha)$. (We set $\alpha=0.3$ in all our experiments.) The eCLIP image encoder then processes $I_{i}^{\lambda}$ to produce the image embedding $v_{i}^{\lambda}=f(I_{i}^{\lambda})$. These expert embeddings form new positive pairs $(v_{i}^{\lambda},t_{i})$ as well as corresponding negative pairs, which are added with existing pairs $(v_{i},t_{i})$ during the computation of the CLIP InfoNCE Loss, $\mathcal{L}_{i}$.
To seamlessly integrate expert annotations without disrupting the foundational training of the eCLIP model, we employ a phased curriculum learning strategy (Bengio et al., 2009). This approach comprises a cold start phase where the model is initially trained without the expert annotations to establish a robust baseline. This phase accounts for about 10% of the total training iterations. It then transitions into a warmup phase, gradually increasing the inclusion of expert examples from 0.05 to 0.5 probability over the next 30% of iterations. Finally, a cooldown phase reduces expert example probability to 0.1 for the subsequent 40% of iterations, fine-tuning the model’s performance by balancing foundational and expert-driven insights.
@@ 4 Experiments
Our experiments are designed to evaluate the influence of expert heatmap annotations on the quality of the learned representations. Unless stated otherwise, we assume that a large set of image-text pairs, of which a small fraction is annotated with eye-gaze heatmaps, is used for training, but for test samples no annotations are used. We utilize both quantitative measures and qualitative assessments to study the contributions of these annotations towards enhancing model performance.
@@ 4.1.1 Baselines
[…]
@@ 4.1.2 Datasets
For pretraining phase we utilize the MIMIC-CXR dataset (Johnson et al., 2019b), which pairs roughly 200K chest X-rays with free-text radiology reports. The images were processed into JPEG format as described in (Johnson et al., 2019a), and the accompanying reports were stripped of unnecessary punctuation and tokenized using the Wordpiece scheme (Devlin et al., 2018). We obtain the eye-gaze heatmap from the EGD-CXR dataset (Demner-Fushman et al., 2016) and process the eye-tracking data to obtain the normalized eye-gaze heatmap which are available for 1080 datapoints.
Our evaluation setup includes multiple publicly available chest X-ray datasets, specifically CheXpert (Irvin et al., 2019), RSNA Pneumonia (Shih et al., 2019), NIH CXR (Wang et al., 2017) and Open-I (Demner-Fushman et al., 2016), each offering a distinct set of imaging and reporting characteristics. Following previous works (Wang et al., 2022; Huang et al., 2021; You et al., 2023), we prepare the test sets from MIMIC and CheXpert, selecting 200 random samples for five specific pathologies from the CheXpert competition, resulting in 1000 samples for each dataset (MIMIC 5x200 and CheXpert 5x200, respectively). For the NIH-CXR dataset, we assembled a subset of 100 samples for each of 14 abnormalities, thereby creating CXR 14x100 test set. The Open-I dataset is utilized for text retrieval and radiology report generation tasks. For linear probe evaluations, we use CXR-8 (Wang et al., 2017), RSNA dataset and construct an OpenI-5 dataset by extracting labels from the ‘Problems’ field within the reports that match the CheXpert competition labels.

### v3

@@ Abstract
Keywords:
Contrastive Learning Medical Imaging Zero-shot Inference
@@ 1 Introduction
Pretraining foundation models on multi-modal data – particularly leveraging the relationships between text and images – has proven to be a robust strategy for generating versatile embeddings [36, 18]. These embeddings enhance the efficacy in several downstream tasks, from image generation to advanced vision-language integration [26, 37, 40]. Central to this approach is the employment of a contrastive learning (CL) loss objective [34, 4, 61], where models are trained to align positive pairs (e.g., an image and its corresponding caption) while diversifying negative ones. A significant hurdle in this approach is the necessity of vast datasets, often comprising several millions of data points, for competitive results. Models such as CLIP [36] have been trained on internet-scale datasets, estimated to encompass hundreds of millions of image-text pairs [5, 57]. Acquiring datasets of this magnitude poses substantial challenges in specialized fields that require expert knowledge for data collection, processing and annotation. The medical imaging domain exemplifies these difficulties, where acquiring even a single data point, such as a chest X-ray, involves complex processes requiring expertise and significant resources. Moreover, the procurement of such data for machine learning research is further complicated by ethical considerations, patient privacy concerns and the need for extensive de-identification procedures.
@@
This has led to the prevalent use of foundation models, initialized with weights from models trained on extensive internet-scale datasets, for tasks in the medical domain [65, 54, 15, 23]. However, the areas of interest within medical images are often nuanced and require expert knowledge to interpret, rendering them indistinguishable to a general-purpose model. In Fig. 1(a), we investigate the embeddings generated by a CLIP model – initially pretrained on internet data – using samples from the Open-I dataset[6], which includes X-rays and corresponding radiology reports. We categorize the samples into subgroups based on the primary abnormality identified in each report, such as ‘normal’, cardiomegaly, atelectasis and opacity. A histogram of the cosine similarities between embeddings from different groups indicates a high degree of similarity, with values approaching 1. This could lead to potential challenges in downstream zero-shot inference tasks, which rely on the spatial segregation of embeddings from different groups [36]. Typically, continual pretraining on medically relevant data is employed to enhance the model’s ability to differentiate between various abnormalities.
@@
(a) Histogram of Cosine Similarities Among Subgroups. The model exhibits high cosine similarity among embeddings within the text (red) and image (blue) modalities, regardless of the differences among subgroups. This underscores the model’s challenge in capturing the subtleties inherent in medical data.
@@
(b) Modality Gap. Despite the CLIP contrastive loss aiming at closely aligning image and text embeddings within a shared space, the modalities remain segregated into distinct regions.
@@
Recent studies [25, 11, 33, 67, 46] have identified a “modality gap” in multi-modal contrastive representation learning, where the embeddings from different modalities (e.g., images and text) fall in distinct regions in the shared embedding space. This separation, which arises from factors such as initial model weights and the objectives of contrastive learning [25, 67], leads to the “cone effect” where embeddings of each modality are restricted to a narrow region of the embedding hypersphere. In Fig. 1(b), we illustrate this within the medical domain with a 2D UMAP[27] projection of image and text embeddings. This example highlights how embeddings from the same modality but different semantic groups, such as X-ray images of varying abnormalities, cluster closely together. This makes it difficult for a model to distinguish between semantically different images, undermining its performance in medical image analysis.
@@
We investigate the potential of integrating expert annotations, specifically radiologist eye-gaze heatmaps, to alleviate these issues. Processing the eye-gaze data from radiologists[22] provides us with heatmaps indicative of the radiologist’s attention across different regions of the X-ray images. This heatmap reflects areas of clinical interest aligned with details present in radiology reports. We posit that this could help capture nuanced visual cues in the X-rays and therefore pairing it with reports can enrich the CLIP training data with high-quality positive pairs. Due to the scarcity of such expert annotated data, we employ the mixup strategy, a data augmentation technique which has been effective in both supervised [62, 48, 13] and contrastive learning[49, 33], to create additional synthetic samples.
@@ 2 Related Work
Modality Gap: Liang et al. [25] pinpoint the origins of the modality gap to the nuances of model initialization and the objectives of contrastive learning, underscoring its impact on downstream tasks and fairness. Oh et al. [33] highlight poor uniformity and alignment in CLIP’s embeddings and propose a finetuning method for robust representations. Zhang et al. [67] explore the geometry of this embedding space, and provide both theoretical and empirical insights on the nature of this geometry. Subsequent research has produced methods to mitigate the modality gap through diverse and creative approaches [11, 46, 66, 12, 32].
Improving Contrastive Learning: Several methods have improved upon the CLIP objective by introducing auxilliary losses, e.g., SLIP [30] uses SimCLR [4] loss, M3MAE [10, 55] augment the Masked Autoencoder [14] reconstruction loss, FLIP [24] randomly masks out input images to improve scaling, DACL [49] proposes a domain agnostic mixup strategy, SILC [31] uses self-distillation, Mo et al. [28] utilized specialist captions to generate pseudo labels for unpaired images, Zhang et al. [63] propose Multi-task Paired Masking with Alignment to improve cross-modal interaction. Similarly there have been works that have identified the need to make the CLIP model focus on sub-regions in order to enhance its utility and downstream performance, GLoRIA [15] considers the loss from local regions from within the image and reports, Alpha-CLIP [45] uses the alpha channel to guide the CLIP model to focus on different regions of the image and generate the masks for all the images in the corpus using an image segmentation pipeline and TIER [35] uses a regularization term to improve the local focus of the model.
Multi-modal Contrastive Learning in Medical Imaging: Zhang et al. [65] demonstrated enhanced downstream performance by jointly using chest X-ray and report pairing for training a contrastive learning model. This was further improved by Huang et al. [15], by exploiting local and global features from both modalities; Wang et al. [54] and You et al. [59] achieved impressive results by using a Swin Tiny model as the image encoder and by adding modifications to the contrastive loss. Several other works have developed similar contrastive learning foundation models while utilizing the biomedical image texts to achieve impressive results [16, 58, 29, 56, 44, 47]. Karargyris et al. [22] and Bigolin et al. [3] augment a subset of the MIMIC-CXR [21] samples with high quality eye-tracking and verbal transcripts from several radiologists. van Sonsbeek et al. [43] and Wang et al. [50] utilize the heatmaps from eye-gaze to improve image classification.
@@ 3.1 Background
Central to CLIP’s effectiveness is the InfoNCE loss [34], a mechanism engineered to optimize the similarity measures between corresponding (positive) pairs and to minimize those among non-corresponding (negative) pairs. The formulation of the CLIP loss objective is as follows:
@@
Theoretical results on CL indicate the concepts of alignment and uniformity as critical for the quality of embeddings [52, 51]. Alignment focuses on reducing the distance between positive pairs while uniformity seeks to evenly distribute the embeddings across the unit hypersphere, preventing extreme clustering that could impair the model’s generalizability and discriminative capabilities. The alignment and uniformity can be defined formally as follows [33]:
@@
Figure 3: Comparing eCLIP with $m^{2}$-mixup[33]. (left) Standard CLIP showing image-text positive pairs $(v_{i},t_{i})$ (solid line), while the other image embeddings serve as negative pairs (dashed line). (center) the $m^{2}$-mixup creates negative pairs $(v^{\lambda}_{j},t_{i})$ via interpolation between embeddings along the geodesic. (right) eCLIP adds expert image embedding, $v^{E}_{i}$, in addition to $v_{i}$ for text $t_{i}$, forming additional positive and negative pairs
High intra-modal similarity, e.g. between two images as seen in Fig. 1, can inadvertently enhance similarity among negative pairs, inflating the denominator of the loss function in Equation (1), and, consequently, hurting the model’s ability to differentiate between positive and negative pairs during training. A conventional method to counter this involves incorporating hard negative pairs, a strategy Oh et al. [33] employ by mixing embeddings from different modalities. While effective, this cross-modal mixup may obscure the semantic clarity of embeddings. As an alternative we propose increasing the dataset with additional positive pairs that exhibit minimal semantic overlap by integrating expert annotations. Fig. 3 compares the positive and negative pair creation in eCLIP with traditional CLIP and $m^{2}$-mixup.
@@ 3.2 Introducing Expert Annotations to CLIP
8:        // Process expert image
9:        $\lambda\sim\text{Beta}(\alpha,\alpha)$
10:        $I_{i}^{E}\leftarrow\text{HeatmapProcessor}(I_{i},E_{i})$
11:        $I_{i}^{\lambda}\leftarrow I_{i}\lambda+I_{i}^{E}(1-\lambda)$
12:        $v_{i}^{\lambda}\leftarrow f(I_{i}^{\lambda})$
13:     end if
14:     if $E_{i}$ is entirely ones then
15:        $I_{i}^{\text{R}}\leftarrow\text{HeatmapProcessor}(I_{i},E_{i})$
16:        $\mathcal{L}_{\text{priming}}\leftarrow\mathcal{L}_{\text{priming}}+(I_{i}-I_{i}^{\text{R}})^{2}$
17:        $n_{\text{p}}\leftarrow n_{\text{p}}+1$
18:     end if
19:  end for
@@
However, the size of the expert annotated data is orders of magnitude smaller than the data available for CLIP training. To effectively leverage the scarce expert-annotated data, we implement mixup augmentation [62]. As illustrated in Fig. 2 (left), this involves blending an original image $I_{i}$ with its expert version $I_{i}^{E}$ to create $I_{i}^{\lambda}=\lambda I_{i}+(1-\lambda)I_{i}^{E}$, where $\lambda\sim\text{Beta}(\alpha,\alpha)$. (We set $\alpha=0.3$ in all our experiments.) The eCLIP image encoder then processes $I_{i}^{\lambda}$ to produce the image embedding $v_{i}^{\lambda}=f(I_{i}^{\lambda})$. These expert embeddings form new positive pairs $(v_{i}^{\lambda},t_{i})$ as well as corresponding negative pairs, which are added with existing pairs $(v_{i},t_{i})$ during the computation of the CLIP InfoNCE Loss, $\mathcal{L}_{i}$.
To seamlessly integrate expert annotations without disrupting the foundational training of the eCLIP model, we employ a phased curriculum learning strategy [2]. This approach comprises a cold start phase where the model is initially trained without the expert annotations to establish a robust baseline. This phase accounts for about 10% of the total training iterations. It then transitions into a warmup phase, gradually increasing the inclusion of expert examples from 0.05 to 0.5 probability over the next 30% of iterations. Finally, a cooldown phase reduces expert example probability to 0.1 for the subsequent 40% of iterations, fine-tuning the model’s performance by balancing foundational and expert-driven insights.
@@ 4 Experiments
Our experiments are designed to evaluate the influence of expert heatmap annotations on the quality of the learned representations. Unless stated otherwise, we assume that a large set of image-text pairs, of which a small fraction is annotated with eye-gaze heatmaps, is used for training, but for test samples no annotations are used. We utilize both quantitative measures and qualitative assessments to study the contributions of these annotations towards enhancing model performance. The source code is available online.
@@ 4.1.1 Baselines
To validate our approach, we compare eCLIP against a model trained using traditional CLIP (referred to as the base model) and a “naive” baseline, where the expert annotated samples are directly added to the training set without using mixup or curriculum learning. We also examine the impact of two mixup methods: Domain Agnostic Contrastive Learning (DACL) [49] and $m^{3}$-mixup [33] which blends image and text embeddings to improve alignment and uniformity across modalities. While DACL is integrated during pretraining, $m^{3}$-mixup is applied post-pretraining, in a manner akin to fine-tuning. eCLIP can be applied to any variant of CLIP, which we demonstrate also with GLoRIA [15] which has a Resnet50 image encoder. We introduce two variants of our technique: eCLIP, which integrates expert annotations during the initial CLIP pretraining phase, and eCLIPP, which instead continually finetunes a trained CLIP model with expert annotations, similar to $m^{3}$-mixup.
@@ 4.1.2 Datasets
For pretraining phase we utilize the MIMIC-CXR dataset [21], which pairs roughly 200K chest X-rays with free-text radiology reports. The images were processed into JPEG format as described in [20], and the accompanying reports were stripped of unnecessary punctuation and tokenized using the Wordpiece scheme [7]. We obtain the eye-gaze heatmap from the EGD-CXR dataset [6] and process the eye-tracking data to obtain the normalized eye-gaze heatmap which are available for 1080 datapoints.
Our evaluation setup includes multiple publicly available chest X-ray datasets, specifically CheXpert [17], RSNA Pneumonia [39], NIH CXR [53] and Open-I [6], each offering a distinct set of imaging and reporting characteristics. Following previous works [54, 15, 59], we prepare the test sets from MIMIC and CheXpert, selecting 200 random samples for five specific pathologies from the CheXpert competition, resulting in 1000 samples for each dataset (MIMIC 5x200 and CheXpert 5x200, respectively). For the NIH-CXR dataset, we assembled a subset of 100 samples for each of 14 abnormalities, thereby creating CXR 14x100 test set. The Open-I dataset is utilized for text retrieval and radiology report generation tasks. For linear probe evaluations, we use CXR-8 [53], RSNA dataset and construct an OpenI-5 dataset by extracting labels from the ‘Problems’ field within the reports that match the CheXpert competition labels.